How many Sulfur(S) atoms are in 15.2 g of Al2 (SO4)3 ?
Answer #1

Mass of Al2(SO4)3 = 15.2 g

Molar mass of Al2(SO4)3 = 342.15 g/mol

Moles of Al2(SO4)3 = Mass/ Molar mass = 15.2 g/ ( 342.15 g/mol) = 0.0444 mol

1 mol of Al2(SO4)3 contains 3 mol of S

so, 0.0444 mol of Al2(SO4)3 contains 3*0.0444 = 0.133 mol of S

1 mol = 6.022 *1023 atoms

So, 0.133 mol = 0.133 mol* 6.022 *1023 atoms / 1mol = 8.03 *1022 atoms